A 10n force acting on a certain mass increases its speed uniformly from 0 to 25 m/s in 5s. The magnitude of the mass in kg is?
1
Expert's answer
2020-04-22T10:03:25-0400

The acceleration of an object

"a=\frac{\Delta v}{\Delta t}=\frac{25}{5}=5\:\rm m/s^2."

The mass of an object

"m=\frac{F}{a}=\frac{10}{5}=2\:\rm kg."